Given the equation 4x-2y+6z=0
What values satisfy the equation when x=2 and z=1?
Define all elements of solution set in which the values of two variables equal zero.
1
Expert's answer
2020-05-03T15:15:12-0400
  1. We substitute the values of the variables x and z into the equation, and we obtain the only value for the variable y: 4*2-2y+6*1=0; 8-2y+6=0; 2y=14; y=7
  2. Let's present our equation in the general form: a1x1+a2x2+a3x3=0. If the values ​​of any two variables are equal to zero, their products by the coefficients of the equation is also equal to zero, and therefore the sum of these products is equal to zero. For example, a1*0+a2*0+a3*x3=0; 0+0+a3x3=0. So we have simple equation a3x3=0. In our case none of the coefficients (a1, a2, a3) of the equation is equal to zero, therefore a3 is not equal to zero. So the only value for x3 is 0. The answer is x=y=z=0.